# EnsemW2S: Can an Ensemble of LLMs be Leveraged to Obtain a Stronger LLM?

### v1

@@ 5.2.2 Comparing Strong model’s performance
Figure 4: Generation Task (Quartz Data): Top figure shows a bar plot comparing the w2s generalization of our method (grey) with a baseline (blue) for various combinations of weak and strong model pairs for the SFT task on Q/A data for random data split (top bar-plot) and easy-hard split (bottom bar-plot). Bottom figure shows a line plot comparing accuracy and PGR. The left two figures are for random data split, while the right two are for the easy-hard split to show e2h generalization.
Note: Refer to Appendix Table 3 and 6 for detailed values of our experiments for Quartz and ARC datasets, respectively, for random data split. Appendix Figure 14 and Fig. 16 show bar plots with weak and strong (oracle) model performance for the Quartz and ARC datasets, respectively, for the random split. For easy-hard data split, the same details can be found in Appendix Tables 4, 7 and Figure 15 and 17.
5.2.3 Performance on hard data after training on weak vs strong data
Quartz ARC
Model Size Easy Split Hard Split Easy Split Hard Split
pythia-70m 49.11 50.13 21.42 25.26
pythia-160m 48.47 46.43 21.85 22.10
pythia-410m 51.50 51.50 18.01 18.95
pythia-1b 53.32 56.77 19.80 22.10
pythia-1.4b 60.34 63.78 21.42 21.42
pythia-2.8b 66.84 70.41 25.09 26.71
Figure 6: Accuracy (%) values for LLMs trained on easy vs hard data and evaluated on hard data.
We conduct this experiment to motivate the importance of e2h with w2s generalization. For the Quartz dataset in Table 6, we see a significant margin of improvement when trained on hard data for the larger models, showing larger models are more capable of understanding complicated data. With ARC, we see improvement in all models but with a lesser margin, implying that ARC data has a lesser disparity between easy and hard samples.
6 Conclusion, Limitation and Future Work
Conclusion: This paper aims to stimulate discussion on the more holistic problem of weak-to-strong generalization by emphasizing easy-to-hard generalization. We develop a new AdaBoost-inspired algorithm and conduct a thought experiment on how to combine the "wisdom of the crowd" to improve w2s generalization. We are first to focus on the idea of making the weaks less weak using an ensemble, and test our method for binary classification and Q/A-based SFT task. Our method in some cases recovers full strong model capability.
Limitation and Future Work: This work only explores the supervised fine-tuning phase. While SFT is an important part of the LLM learning pipeline, our future work will focus on developing weak supervision in the reward modeling phase. Another interesting future direction would be to improve the combination of tokens in the decoding phase by replacing the classical AdaBoost algorithm with more adaptive ensemble learning methods. We hope this work sparks discussion on combining multiple LLMs to improve weak-to-strong generalization.
Acknowledgements
@@
Agrawal, Ding, Che, Deng, Satheesh, Langford and Huang are supported by DARPA Transfer from Imprecise and Abstract Models to Autonomous Technologies (TIAMAT) 80321, National Science Foundation NSF-IIS-2147276 FAI, DOD-ONR-Office of Naval Research under award number N00014-22-1-2335, DOD-AFOSR-Air Force Office of Scientific Research under award number FA9550-23-1-0048, DOD-DARPA-Defense Advanced Research Projects Agency Guaranteeing AI Robustness against Deception (GARD) HR00112020007, Adobe, Capital One and JP Morgan faculty fellowships.
Appendix A Related Works
@@
Additionally, Mudgal et al. (2023) introduces Controlled Decoding (CD), a method for aligning language model outputs with desired outcomes using a separate prefix scorer module. This approach allows multi-objective RL without additional training and performs well on benchmarks, bridging the gap between token-level control and sequence-level best-of sampling strategies.
Appendix B Limitation and Future Work
(Continue from main manuscript)
Appendix C Detailed Flowchart

### v2

@@ 6.1 Enhancing ID and OOD Generalization of Weak Experts using EnsemW2S
Figure 4: (Quartz Data.) Top: Bar plots comparing W2S generalization of the strong student using our method (gray) vs. baseline (blue) across weak–strong model pairs for SFT on Q/A data—ID (top bars) and OOD (bottom bars). Bottom: Line plots showing accuracy and PGR. Left: ID (random split); Right: OOD (easy–hard split) for E2H generalization.
Following Burns et al. (2023), we study scaling laws for W2S generalization in Fig.4 by creating multiple weak–strong model pairs from the same model family. We use the Pythia model series, which offers a wide range of model sizes. To enable a direct comparison with Burns et al. (2023), we train the strong model using only hard data. Results for the Quartz dataset are shown here, while results for ARC are included in AppendixJ.2. We also conduct experiments with a binary classification LLM on the SciQ task using AdaBoost as the ensemble method, following the exact experimental setup and code provided by Burns et al. (2023) (Results present in Appendix J.3). For all above experiments, we observe consistent improvements—indicating stronger W2S generalization—except in cases where models are too small to learn meaningful information. In the binary setting, we even observe instances of complete W2S generalization, demonstrating the effectiveness of carefully constructed weak experts and highlighting the overall merit of the ensemble approach.
7 Conclusion
This paper presents a method to iteratively train weak experts by exposing them only to human-level labeled data, enabling them to better generalize to superhuman-level data. These weak supervisors are then combined at the token level to generate stronger supervision, improving W2S generalization for the strong student model. Our method shows significant improvement on both in-distribution and out-of-distribution datasets.
8 Acknowledgement
@@
Agrawal, Ding, Deng, Che, Satheesh, An and Huang are supported by DARPA Transfer from Imprecise and Abstract Models to Autonomous Technologies (TIAMAT) 80321, National Science Foundation NSF-IIS-2147276 FAI, DOD-AFOSR-Air Force Office of Scientific Research under award number FA9550-23-1-0048, Adobe, Capital One and JP Morgan faculty fellowships.
Appendix A Limitation and Future Work
This work only explores the SFT phase. While SFT is an important part of the LLM learning pipeline, our future work will focus on developing weak supervision in the reward modeling phase. Another interesting future direction would be to improve the combination of tokens in the decoding phase by replacing the classical AdaBoost algorithm with more adaptive ensemble learning methods. We hope this work sparks discussion on combining multiple LLMs to improve w2s generalization.
Appendix B Broader Impact
Appendix C Related Works
@@
Additionally, Mudgal et al. (2023a) introduces Controlled Decoding (CD), a method for aligning language model outputs with desired outcomes using a separate prefix scorer module. This approach allows multi-objective RL without additional training and performs well on benchmarks, bridging the gap between token-level control and sequence-level best-of sampling strategies.
Appendix D Adaboost for combing weak Binary classification LLM.
AdaBoost is an ensemble learning algorithm that combines multiple weak classifiers, such as decision stumps, to create a strong classifier. It works iteratively by focusing on the samples that are hardest to classify, assigning them higher weights in each subsequent iteration. Weak classifiers are trained one at a time, and their contributions are weighted based on their accuracy. The final prediction is made by taking a weighted majority vote of all weak classifiers. AdaBoost is known for its ability to improve generalization by focusing on difficult cases and is often resistant to overfitting with simple weak learners. However, it can struggle with noisy data if overemphasis is placed on misclassified samples. Its also presented as Algorithm 2.
We use Adaboost to conduct a thought experiment to test the ensemble idea in w2s generalization for simple classification task. This is also the first task evaluated by Burns et al. (2023). We use vanilla AdaBoost (Algorithm 2) to generate answers to hard questions, $\bm{x}^{h}$, from each weak LLM teacher, $h_{\theta}^{t}(\bm{x}^{h})$ for $t\in\{1,\ldots,T\}$, where $T$ is max Adaboost round. AdaBoost iteratively reweights training examples, and trains weak teachers one at a time focusing on hard-to-classify samples, as described in Line 5 of Algorithm 2. The only requirement is they perform better than random, thus satisfying the weak learning condition. A weighted majority vote aggregates their outputs to produce a pseudo-label, $\mathbbm{1}(\sum_{t=1}^{T}\alpha_{t}h_{\theta}^{t}(\bm{x}^{h})>0)\in\{0,1\}$, where ${\alpha_{t}}$ are hyperparameters weighing each weak teacher based on accuracy. A detailed mathematical summary is provided below.
Let the training dataset consist of $m$ samples:
$\{(x_{i},y_{i})\mid i=1,2,\dots,m\},\quad x_{i}\in\mathbb{R}^{d},\quad y_{i}\in\{-1,+1\}.$
Each weak learner $h_{t}(x)$ outputs a prediction $h_{t}(x_{i})\in\{-1,+1\}$. The goal is to sequentially train weak learners such that the combined model minimizes the classification error. A weight distribution $D_{t}(i)$ is maintained over the training samples at each iteration $t$, where:
$D_{t}(i)\geq 0,\quad\sum_{i=1}^{m}D_{t}(i)=1.$
Initially, all samples are equally weighted: $D_{1}(i)=\frac{1}{m},\quad\forall i$
Training the Weak Learners:
For each iteration $t=1,2,\dots,T$, train a weak learner $h_{t}(x)$ using the current weight distribution $D_{t}$. Compute the weighted error:
$\epsilon_{t}=\sum_{i=1}^{m}D_{t}(i)\cdot\mathbb{I}(h_{t}(x_{i})\neq y_{i}),$
where $\mathbb{I}(\cdot)$ is the indicator function.
Weak Learner Weight
Assign a weight $\alpha_{t}$ to the weak learner based on its performance:
$\alpha_{t}=\frac{1}{2}\ln\left(\frac{1-\epsilon_{t}}{\epsilon_{t}}\right)$
Intuition behind it is that if $\epsilon_{t}$ is small, $\alpha_{t}$ is large, giving more importance to the weak learner. If $\epsilon_{t}=0.5$, $\alpha_{t}=0$, indicating no contribution to the ensemble. $\epsilon_{t}>0.5$ is undesirable, as the weak learner performs worse than random guessing.
Update the weights of the training samples to focus on misclassified samples:
$D_{t+1}(i)=\frac{D_{t}(i)\exp(-\alpha_{t}y_{i}h_{t}(x_{i}))}{Z_{t}},$
where $Z_{t}$ is a normalization factor ensuring $\sum_{i=1}^{m}D_{t+1}(i)=1$:
$Z_{t}=\sum_{i=1}^{N}D_{t}(i)\exp(-\alpha_{t}y_{i}h_{t}(x_{i})).$
Misclassified samples ($y_{i}\neq h_{t}(x_{i})$) receive higher weights, making them more influential in the next iteration. The final strong classifier $H(x)$ is a weighted majority vote of the weak learners:
$H(x)=\operatorname{sign}\left(\sum_{t=1}^{T}\alpha_{t}h_{t}(x)\right).$
Generalization Abilities:
AdaBoost improves generalization by maximizing the margins on the training set. The margin for a sample $(x_{i},y_{i})$ is defined as:
$\text{Margin}(x_{i})=y_{i}\sum_{t=1}^{T}\alpha_{t}h_{t}(x_{i}).$
AdaBoost aims to increase the margin for all samples, reducing the chance of misclassification.
Summary of Key Properties
  1. a)
    Sequential Training: Weak learners are trained iteratively, with weights updated to focus on difficult samples.
  2. b)
    Weighting Scheme: Misclassified samples are emphasized in subsequent iterations.
  3. c)
    Generalization: AdaBoost achieves strong generalization by maximizing margins and minimizing exponential loss.
  4. d)
    Flexibility: It can work with any weak learner as long as the learner achieves performance slightly better than random guessing.
Algorithm 2 AdaBoost Freund & Schapire (1997)
0:  Training Dataset $S=\{(x_{i},y_{i})\}_{i=1}^{m}\sim D^{m}$
  $T=$ AdaBoost iterations
  $\vec{D}_{1}(i)\leftarrow\frac{1}{m}\forall i\in[m]$
  for $t\leftarrow 1$ to $T$  do
     $\alpha_{t}\leftarrow\frac{1}{2}\log{\frac{1-\epsilon_{t}}{\epsilon_{t}}}$
     $Z_{t}\leftarrow 2\sqrt{\epsilon_{t}(1-\epsilon_{t})}$
     $\vec{D}_{t+1}\leftarrow\frac{1}{Z_{t}}\vec{D}_{t}e^{-\alpha_{t}y_{i}h_{t}(x_{i})}$
     $g\leftarrow\sum_{t=1}^{T}\alpha_{t}h_{t}$
  Return $h(x)=\operatorname{sign}(g)$
Appendix E Details on the EnemW2S Methodology
E.1 Detailed Flowchart